# Refining Salience-Aware Sparse Fine-Tuning Strategies for Language Models

### v1

@@ Abstract
Parameter-Efficient Fine-Tuning (PEFT) has gained prominence through low-rank adaptation methods like LoRA. In this paper, we focus on sparsity-based PEFT (SPEFT), which introduces trainable sparse adaptations to the weight matrices in the model, offering greater flexibility in selecting fine-tuned parameters compared to low-rank methods. We conduct the first systematic evaluation of salience metrics for SPEFT, inspired by zero-cost NAS proxies, and identify simple gradient-based metrics is reliable, and results are on par with the best alternatives, offering both computational efficiency and robust performance. Additionally, we compare static and dynamic masking strategies, finding that static masking, which predetermines non-zero entries before training, delivers efficiency without sacrificing performance, while dynamic masking offers no substantial benefits. Across NLP tasks, a simple gradient-based, static SPEFT consistently outperforms other fine-tuning methods for LLMs, providing a simple yet effective baseline for SPEFT. Our work challenges the notion that complexity is necessary for effective PEFT, and is open-source and available to the community [^1] Available at: https://github.com/0-ml/speft. .
@@ 1 Introduction
Pretrained large language models (LLMs) have demonstrated strong performance across various natural language processing (NLP) tasks [4]. A typical approach for adapting these LLMs to specific downstream tasks involves fine-tuning their trainable parameters. However, this process can be prohibitively expensive on consumer-grade hardwares, if we consider training all free parameters, especially on LLMs exceeding a billion parameters. For example, models with over 100 billion parameters, such as BLOOM, required training with 384 GPUs across 48 distributed computing nodes [24]. Instead of training all parameters, an alternative fine-tuning paradigm that enables model training on new tasks with minimal computational resources is Parameter-Efficient Fine-Tuning (PEFT). This method aims to learn only a small set of parameters in order to adapt the model to the new task, substantially lowers the computational resource requirements [2, 15].
Figure 1: Comparison between LoRA [15] and SPEFT. LoRA freezes pretrained weights ${\bm{\theta}}_{0}$ and updates the low-rank terms $A$ and $B$, while SPEFT adopts zero-cost proxies to build a sparse adapter ${\bm{\theta}}_{\mathrm{sp}}$, to update the weight elements that contribute most to the downstream task.
Existing effort on PEFT methods mainly focuses on two categorizes, low-rank-based and sparsity-based adaptation approaches. LoRA [15], a popular low-rank adaptation method reparameterizes the model weight of each layer (${\bm{\theta}}\in\mathbb{R}^{d_{1}\times d_{2}}$) as ${\bm{\theta}}\triangleq{\bm{\theta}}_{0}+BA$, where ${\bm{\theta}}_{0}$ denotes the pretrained weight matrix which remains fixed during fine-tuning, $B\in\mathbb{R}^{d_{1}\times r}$ and $A\in\mathbb{R}^{r\times d_{2}}$ are trainable weights of a lower rank with $r\ll\min\{d_{1},d_{2}\}$. Recently, sparsity-based PEFT (SPEFT) has emerged as an alternative approach which constructs an alternate reparameterization, ${\bm{\theta}}\triangleq{\bm{\theta}}_{0}+{\bm{\theta}}_{\mathrm{sp}}$, where ${\bm{\theta}}_{\mathrm{sp}}$ is an extremely sparse matrix, and updates solely its non-zero entries. Figure 1 illustrates the distinction between the two categories of PEFT methods. Previous sparse PEFT methods [12, 30, 2] have employed various first- and second-order metrics for determining these non-zero entries and adopted distinct approaches for handling the sparsity mask during training. The varying constructions and training-time treatments of the sparsity mask lead us to the following research questions on the basic design principles for SPEFT:
@@
In this paper, we systematically re-examine the design principles for SPEFT and conduct an evaluation across distinct salience metrics. Drawing inspiration from recent advancements in zero-cost Network Architecture Search (NAS) proxies, which explore diverse low-cost proxies for determining parameter importance that has incorporated both first-order (e.g.missing, weight magnitude, gradients, SNIP [21], etc.missing) and second-order estimators (e.g.missing, GRaSP [35], Fisher information [30], etc.missing), we discovered that these NAS proxies encompasses many salience metrics used in SPEFT for sparsity mask construction (DiffPruning [12], FishMASK [30], etc.missing). Consequently, inspired by recent zero-cost NAS metrics that have shown strong performance to construct sparsity masks, we are the first to comprehensively evaluate 8 different salience metrics in the context of SPEFT for LLMs. Furthermore, we investigate both dynamic and static masking approaches, where a dynamic mask matrix ${\bm{\tau}}$ changes during training, while a static mask maintains a static ${\bm{\tau}}$ binary matrix throughout the PEFT process. We make the following contributions:
@@
  • Our results indicate that a simple gradient-based, static SPEFT method delivers the best trade-off between effectiveness and efficiency. For instance, for RoBERTa-base on MRPC task, our method achieves 0.98% higher than the baseline given the same amount of trainable parameters. Gradient-based SPEFT outperforms LoRA by 22.6% on GSM8k [7] when trained on MetaMathQA [38]. Consequently, we advocate for this SPEFT variant to be considered a strong baseline for subsequent developments in this field.
@@ 2.1 PEFT Methods
Low-rank adaptation [15] is a popular method in PEFT which reparameterizes the weight matrix of each layer (${\bm{\theta}}\in\mathbb{R}^{d_{1}\times d_{2}}$) as ${\bm{\theta}}={\bm{\theta}}_{0}+BA$. Here, ${\bm{\theta}}_{0}\in\mathbb{R}^{d_{1}\times d_{2}}$ is the pretrained weight matrix, and $B\in\mathbb{R}^{d_{1}\times r}$ and $A\in\mathbb{R}^{r\times d_{2}}$ are lower-rank matrices with $r\ll\min(d_{1},d_{2})$. By making only $A$ and $B$ trainable, this method significantly reduces the number of trainable parameters, thereby lowering computational resource requirements. LoRA has demonstrated effectiveness in reducing trainable parameters for fine-tuning large language models, while maintaining strong fine-tuned performance across various downstream tasks compared to full fine-tuning.
@@
This approach has been explored in various forms in the literature. Earlier works such as DiffPruning [12] learns a sparsity mask with straight-through gradient estimator [3, 16] to select important parameters for downstream tasks. FishMASK [30] applies a static sparsity mask from training outset, guided by Fisher information to measure sparsity. Beyond static masks, Fish-DIP [8] further allows the Fisher information-based mask to be updated dynamically during training. Inspired by the lottery ticket hypothesis [11], LF-SFT [2] finds that sparse masks obtained by selecting the parameters with the largest changes after fine-tuning on a task can be transferred to other tasks. However, this approach requires full fine-tuning on an initial task, which may not be feasible for resource-constrained settings. This paper explores the design principles for constructing the sparsity mask with low-cost salience metrics and the impact of dynamic versus static masks on the fine-tuning process.
Finally, sparsity-based adapters also allows highly granular control over trainable parameters, and can enable the use of existing knowledge transfer techniques, such as mixtures of sparse experts [36] and multi-task learning with sparse masks [29] in LLMs.
@@ 2.2 Salience Proxies for Sparsity Masking
The extensive research on low-cost salience metrics for fine-grained network pruning has provided a rich set of pruning-at-initialization metrics to determine the importance of neural network parameters. These metrics can be broadly classified into first- and second-order categories. First-order metrics include weight magnitude [13], connection sensitivity (SNIP) [20], foresight connection sensitivity (FORCE) [9], Taylor-FO [26], SynFlow [31], and finally, the gradient of the loss with respect to the weight. Second-order metrics comprise GRaSP [35] and Fisher information-based metrics [22]. Coincidentally, both FishMASK [30] and Fish-DIP [8] propose to use Fisher information to construct the sparsity mask: while FishMASK uses a static mask, Fish-DIP further allows the mask to be updated periodically during fine-tuning. These metrics are designed to identify important parameters or connections in a neural network. In this paper, we explore the impact of these salience metrics on fine-tuning by using them to construct sparse masks for PEFT.
@@ 3.2 Salience Metrics
  • SNIP (single-shot network pruning): $\left\lvert\frac{\partial{\ell}}{\partial{{\bm{\theta}}}}\odot{\bm{\theta}}\right\rvert$, the connection sensitivity metric proposed in [20] to determine the importance of weights.
  • FORCE (foresight connection sensitivity): $-\frac{\partial{\ell}}{\partial{{\bm{\theta}}}}\odot{\bm{\theta}}$, introduced in [9].
  • Taylor-FO (Taylor first-order expansion): ($\left\lparen\frac{\partial{\ell}}{\partial{{\bm{\theta}}}}\odot{\bm{\theta}}\right\rparen^{2}$, derived from the 1st-order Taylor expansion of the loss [26].
  • SynFlow (iterative synaptic flow pruning): $\frac{\partial{}}{\partial{{\bm{\theta}}}}\left[\mathbf{1}^{\top}\left\lparen\Pi_{l=1}^{L}\lvert{\bm{\theta}}^{(l)}\rvert\right\rparen\mathbf{1}\right]\odot{\bm{\theta}}$, where ${\bm{\theta}}^{(l)}$ denotes the weights of the $l$th layer, and $L$ denotes the number of layers. A data-free metric proposed in [31] to model synaptic flow.
@@
  • GRaSP (gradient signal preservation): ($-\left\lparen H\frac{\partial{\ell}}{\partial{{\bm{\theta}}}}\right\rparen\odot{\bm{\theta}}$, which is a 2nd-order metric proposed in [35] that aims to preserve gradient signals rather than the loss value.
  • Fisher information: ($\left\lparen\frac{\partial{\ell}}{\partial{{\bm{\theta}}}}\right\rparen^{2}$, which uses the Fisher information to determine the importance of weights [30, 8].
@@ 4 Experimental Results
We evaluated our approaches and baselines over a set of models, including fine-tuned OPT variants (-125m, -350m, and -1.3b) [39], BERT-base-uncased [10] and RoBERTa-base [23], for the GLUE [34] benchmark, and fine-tuned Gemma2-2b [33] and Qwen2-7b [37], to evaluate on the Massive Multitask Language Understanding (MMLU) benchmark [14] and GSM8K [7], a dataset of grade school math problems. In addition to sparse PEFT methods presented in this paper, we further include LoRA [15] and PiSSA [25] as low-rank adapter baselines for comparison.
@@
To show the generality of our approach, we chose GLUE, MMLU and GSM8K as benchmarks for evaluation. For the GLUE [34] benchmark, six representative tasks with large sizes are selected: single-sentence task SST-2, inference tasks QNLI, MNLI, similarity and paraphrase tasks MRPC, STS-B and QQP [^2] We did not evaluate CoLA and RTE because these datasets are too small and require special treatments such as fine-tuning RTE using an MNLI checkpoint [18].. For the MMLU [14] benchmark, it contains questions covering 57 subjects across STEM, the humanities, the social sciences, and others. It is designed to test the model’s ability to handle various types of language data and complex problems. We fine-tuned Gemma-2-2b, Qwen2-7b on either the Alpaca [32] or OASST2 [17] conversational datasets, and then evaluated them on all tasks in MMLU. We fine-tuned Gemma2-2b on the MetaMathQA [38] dataset and evaluated on GSM8K (8-shots) to assess the models’ multi-step mathematical reasoning ability. In the results, we reported the match accuracy for MNLI, Pearson correlation for STS-B, flexible extract and strict match scores for GSM8K, and accuracy values for other tasks.
@@
We chose LoRA [15] and PiSSA [25] as the competing low-rank baselines across models and benchmarks. By default in all comparisons, SPEFT methods use global sparsity ranking with static masks. For statistical significance, we repeated each experiment 3 times for OPT-{125m,350m}, BERT-base-uncased, and RoBERTa-base, and reported average metrics and their standard deviations.
@@
We also used the most reliable salience metric, i.e.missing, gradient-based, in further experiments to explore questions related to dynamic vs.missing static masks, and global vs.missing local sparsity in Section 4.3.

### v2

@@ Abstract
Parameter-Efficient Fine-Tuning (PEFT) has gained prominence through low-rank adaptation methods like LoRA. In this paper, we focus on sparsity-based PEFT (SPEFT), which introduces trainable sparse adaptations to the weight matrices in the model, offering greater flexibility in selecting fine-tuned parameters compared to low-rank methods. We conduct the first systematic evaluation of salience metrics for SPEFT, inspired by zero-cost NAS proxies, and identify simple gradient-based metrics is reliable, and results are on par with the best alternatives, offering both computational efficiency and robust performance. Additionally, we compare static and dynamic masking strategies, finding that static masking, which predetermines non-zero entries before training, delivers efficiency without sacrificing performance, while dynamic masking offers no substantial benefits. Across NLP tasks, a simple gradient-based, static SPEFT consistently outperforms other fine-tuning methods for LLMs, providing a simple yet effective baseline for SPEFT. Our work challenges the notion that complexity is necessary for effective PEFT, while our open-source framework establishes a reproducible benchmark for future research [^1] Available at: https://github.com/0-ml/speft. .
Refining Salience-Aware Sparse Fine-Tuning Strategies for Language Models
Xinxin Liu1,2                        Aaron Thomas3                        Cheng Zhang4
Jianyi Cheng5                        Yiren Zhao4                        Xitong Gao2,6thanks: Corresponding author, xt.gao@siat.ac.cn.
1 Southern University of Science and Technology 2 Shenzhen Institutes of Advanced Technology, CAS 3 University of Birmingham    4 Imperial College London    5 University of Edinburgh 6 Shenzhen University of Advanced Technology
@@ 1 Introduction
Pretrained large language models (LLMs) have demonstrated strong performance across various natural language processing (NLP) tasks Brown et al. (2020). A typical approach for adapting these LLMs to specific downstream tasks involves fine-tuning their trainable parameters. However, this process can be prohibitively expensive on consumer-grade hardwares, if we consider training all free parameters, especially on LLMs exceeding a billion parameters. For example, models with over 100 billion parameters, such as BLOOM, required training with 384 GPUs across 48 distributed computing nodes Luccioni et al. (2023). Instead of training all parameters, an alternative fine-tuning paradigm that enables model training on new tasks with minimal computational resources is Parameter-Efficient Fine-Tuning (PEFT). This method aims to learn only a small set of parameters in order to adapt the model to the new task, substantially lowers the computational resource requirements Ansell et al. (2021); Hu et al. (2021).
Figure 1: Comparison between LoRA Hu et al. (2021) and SPEFT. LoRA freezes pretrained weights ${\bm{\theta}}_{0}$ and updates the low-rank terms $A$ and $B$, while SPEFT adopts zero-cost proxies to build a sparse adapter ${\bm{\theta}}_{\mathrm{sp}}$, to update the weight elements that contribute most to the downstream task.
Existing effort on PEFT methods mainly focuses on two categorizes, low-rank-based and sparsity-based adaptation approaches. LoRA Hu et al. (2021), a popular low-rank adaptation method reparameterizes the model weight of each layer (${\bm{\theta}}\in\mathbb{R}^{d_{1}\times d_{2}}$) as ${\bm{\theta}}\triangleq{\bm{\theta}}_{0}+BA$, where ${\bm{\theta}}_{0}$ denotes the pretrained weight matrix which remains fixed during fine-tuning, $B\in\mathbb{R}^{d_{1}\times r}$ and $A\in\mathbb{R}^{r\times d_{2}}$ are trainable weights of a lower rank with $r\ll\min\{d_{1},d_{2}\}$. Recently, sparsity-based PEFT (SPEFT) has emerged as an alternative approach which constructs an alternate reparameterization, ${\bm{\theta}}\triangleq{\bm{\theta}}_{0}+{\bm{\theta}}_{\mathrm{sp}}$, where ${\bm{\theta}}_{\mathrm{sp}}$ is an extremely sparse matrix, and updates solely its non-zero entries. Figure 1 illustrates the distinction between the two categories of PEFT methods. Previous sparse PEFT methods Guo et al. (2020); Sung et al. (2021); Ansell et al. (2021) have employed various first- and second-order metrics for determining these non-zero entries and adopted distinct approaches for handling the sparsity mask during training. The varying constructions and training-time treatments of the sparsity mask lead us to the following research questions on the basic design principles for SPEFT:
@@
In this paper, we systematically re-examine the design principles for SPEFT and conduct an evaluation across distinct salience metrics. Drawing inspiration from recent advancements in zero-cost Network Architecture Search (NAS) proxies, which explore diverse low-cost proxies for determining parameter importance that has incorporated both first-order (e.g.missing, weight magnitude, gradients, SNIP Lee et al. (2019b), etc.missing) and second-order estimators (e.g.missing, GRaSP Wang et al. (2020), Fisher information Sung et al. (2021), etc.missing), we discovered that these NAS proxies encompasses many salience metrics used in SPEFT for sparsity mask construction (DiffPruning Guo et al. (2020), FishMASK Sung et al. (2021), etc.missing). Consequently, inspired by recent zero-cost NAS metrics that have shown strong performance to construct sparsity masks, we are the first to comprehensively evaluate 8 different salience metrics in the context of SPEFT for LLMs. Furthermore, we investigate both dynamic and static masking approaches, where a dynamic mask matrix ${\bm{\tau}}$ changes during training, while a static mask maintains a static ${\bm{\tau}}$ binary matrix throughout the PEFT process. We make the following contributions:
@@
  • Our results indicate that a simple gradient-based, static SPEFT method delivers the best trade-off between effectiveness and efficiency. For instance, for RoBERTa-base Liu et al. (2019) on MRPC Dolan and Brockett (2005) task, our method achieves 0.98% higher than the baseline given the same amount of trainable parameters. Gradient-based SPEFT outperforms LoRA by 22.6% on GSM8k Cobbe et al. (2021) when trained on MetaMathQA Yu et al. (2024). Consequently, we advocate for this SPEFT variant to be considered a strong baseline for subsequent developments in this field.
@@ 2.1 PEFT Methods
Low-rank adaptation Hu et al. (2021) is a popular method in PEFT which reparameterizes the weight matrix of each layer (${\bm{\theta}}\in\mathbb{R}^{d_{1}\times d_{2}}$) as ${\bm{\theta}}={\bm{\theta}}_{0}+BA$. Here, ${\bm{\theta}}_{0}\in\mathbb{R}^{d_{1}\times d_{2}}$ is the pretrained weight matrix, and $B\in\mathbb{R}^{d_{1}\times r}$ and $A\in\mathbb{R}^{r\times d_{2}}$ are lower-rank matrices with $r\ll\min(d_{1},d_{2})$. By making only $A$ and $B$ trainable, this method significantly reduces the number of trainable parameters, thereby lowering computational resource requirements. LoRA has demonstrated effectiveness in reducing trainable parameters for fine-tuning large language models, while maintaining strong fine-tuned performance across various downstream tasks compared to full fine-tuning.
@@
This approach has been explored in various forms in the literature. Earlier works such as DiffPruning Guo et al. (2020) learns a sparsity mask with straight-through gradient estimator Bengio et al. (2013); Hubara et al. (2016) to select important parameters for downstream tasks. FishMASK Sung et al. (2021) applies a static sparsity mask from training outset, guided by Fisher information to measure sparsity. Beyond static masks, Fish-DIP Das et al. (2023) further allows the Fisher information-based mask to be updated dynamically during training. Inspired by the lottery ticket hypothesis Frankle and Carbin (2019), LF-SFT Ansell et al. (2021) finds that sparse masks obtained by selecting the parameters with the largest changes after fine-tuning on a task can be transferred to other tasks. However, this approach requires full fine-tuning on an initial task, which may not be feasible for resource-constrained settings. This paper explores the design principles for constructing the sparsity mask with low-cost salience metrics and the impact of dynamic versus static masks on the fine-tuning process.
Finally, sparsity-based adapters also allow highly granular control over trainable parameters, and can enable the use of existing knowledge transfer techniques, such as mixtures of sparse experts Xu et al. (2024) and multi-task learning with sparse masks Sun et al. (2020) in LLMs.
@@ 2.2 Salience Proxies for Sparsity Masking
The extensive research on low-cost salience metrics for fine-grained network pruning has provided a rich set of pruning-at-initialization metrics to determine the importance of neural network parameters. These metrics can be broadly classified into first- and second-order categories. First-order metrics include weight magnitude Han et al. (2015), connection sensitivity (SNIP) Lee et al. (2019a), foresight connection sensitivity (FORCE) de Jorge et al. (2021), Taylor-FO Molchanov et al. (2019), SynFlow Tanaka et al. (2020), and finally, the gradient of the loss with respect to the weight. Second-order metrics comprise GRaSP Wang et al. (2020) and Fisher information-based metrics Liu et al. (2021). Coincidentally, both FishMASK Sung et al. (2021) and Fish-DIP Das et al. (2023) propose to use Fisher information to construct the sparsity mask: while FishMASK uses a static mask, Fish-DIP further allows the mask to be updated periodically during fine-tuning. These metrics are designed to identify important parameters or connections in a neural network. In this paper, we explore the impact of these salience metrics on fine-tuning by using them to construct sparse masks for PEFT.
@@ 3.2 Salience Metrics
  • SNIP (single-shot network pruning): $\left\lvert\frac{\partial{\ell}}{\partial{{\bm{\theta}}}}\odot{\bm{\theta}}\right\rvert$, the connection sensitivity metric proposed in Lee et al. (2019a) to determine the importance of weights.
  • FORCE (foresight connection sensitivity): $-\frac{\partial{\ell}}{\partial{{\bm{\theta}}}}\odot{\bm{\theta}}$, introduced in de Jorge et al. (2021).
  • Taylor-FO (Taylor first-order expansion): ($\left\lparen\frac{\partial{\ell}}{\partial{{\bm{\theta}}}}\odot{\bm{\theta}}\right\rparen^{2}$, derived from the 1st-order Taylor expansion of the loss Molchanov et al. (2019).
  • SynFlow (iterative synaptic flow pruning): $\frac{\partial{}}{\partial{{\bm{\theta}}}}\left[\mathbf{1}^{\top}\left\lparen\Pi_{l=1}^{L}\lvert{\bm{\theta}}^{(l)}\rvert\right\rparen\mathbf{1}\right]\odot{\bm{\theta}}$, where ${\bm{\theta}}^{(l)}$ denotes the weights of the $l$th layer, and $L$ denotes the number of layers. A data-free metric proposed in Tanaka et al. (2020) to model synaptic flow.
@@
  • GRaSP (gradient signal preservation): ($-\left\lparen H\frac{\partial{\ell}}{\partial{{\bm{\theta}}}}\right\rparen\odot{\bm{\theta}}$, which is a 2nd-order metric proposed in Wang et al. (2020) that aims to preserve gradient signals rather than the loss value.
  • Fisher information: ($\left\lparen\frac{\partial{\ell}}{\partial{{\bm{\theta}}}}\right\rparen^{2}$, which uses the Fisher information to determine the importance of weights Sung et al. (2021); Das et al. (2023).
@@ 4 Experimental Results
We evaluated our approaches and baselines over a set of models, including fine-tuned OPT variants (-125m, -350m, and -1.3b) Zhang et al. (2022), BERT-base-uncased Devlin et al. (2019) and RoBERTa-base Liu et al. (2019), for the GLUE Wang et al. (2019) benchmark, and fine-tuned Gemma2-2b Team et al. (2024) and Qwen2-7b Yang et al. (2024), to evaluate on the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2021) and GSM8K Cobbe et al. (2021), a dataset of grade school math problems. Moreover, we fine-tuned Llama3-8b Grattafiori et al. (2024) to evaluate on the HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) benchmarks. In addition to sparse PEFT methods presented in this paper, we further include LoRA Hu et al. (2021) and PiSSA Meng et al. (2024) as low-rank adapter baselines for comparison.
@@
To show the generality of our approach, we chose GLUE, MMLU, GSM8K, HumanEval and MBPP as benchmarks for evaluation. For the GLUE Wang et al. (2019) benchmark, six representative tasks with large sizes are selected: single-sentence task SST-2, inference tasks QNLI, MNLI, similarity and paraphrase tasks MRPC, STS-B and QQP [^2] We did not evaluate CoLA and RTE because these datasets are too small and require special treatments such as fine-tuning RTE using an MNLI checkpoint Lan et al. (2019).. For the MMLU Hendrycks et al. (2021) benchmark, it contains questions covering 57 subjects across STEM, the humanities, the social sciences, and others. It is designed to test the model’s ability to handle various types of language data and complex problems. We fine-tuned Gemma-2-2b, Qwen2-7b on either the Alpaca Taori et al. (2023) or OASST2 Köpf et al. (2023) conversational datasets, and then evaluated them on all tasks in MMLU. We fine-tuned Gemma2-2b on the MetaMathQA Yu et al. (2024) dataset and evaluated on GSM8K (5-shots) to assess the models’ multi-step mathematical reasoning ability. Furthermore, we fine-tuned Llama3-8b on the CodeFeedback Zheng et al. (2024), to evaluate on the HumanEval and MBPP which are aimed to test the code generation ability of models. In the results, we reported the match accuracy for MNLI, Pearson correlation for STS-B, flexible extract and strict match scores for GSM8K, Pass@1 for HumanEval and MBPP, and accuracy values for other tasks.
@@
We chose LoRA Hu et al. (2021) and PiSSA Meng et al. (2024) as the competing low-rank baselines across models and benchmarks. By default in all comparisons, SPEFT methods use global sparsity ranking with static masks. For statistical significance, we repeated each experiment 3 times for OPT-{125m,350m}, BERT-base-uncased, and RoBERTa-base, and reported average metrics and their standard deviations.
@@
We used the most reliable salience metric, i.e.missing, gradient-based, in further experiments to explore questions related to dynamic vs.missing static masks, and global vs.missing local sparsity in Section 4.3. Additionally, we also explored the efficiency-performance trade-off between LoRA, PiSSA and sparse baselines in Appendix C.